If v∈R2 lies in the first quadrant and makes an angle π/4 with the positive x-axis and ||v||=2, then, Select one:


a. v=⟨2√2, 2√2⟩


b. v=⟨2, 2√3⟩


c. v=⟨√2, √2⟩


d. v=⟨2, −2√3⟩


e. v=⟨2√2, −√2⟩
1
Expert's answer
2020-05-13T18:59:36-0400

c)"v=(\sqrt{2};\sqrt{2})"

"v" lies in the first quadrant, hence "x\ge 0" and "y\ge 0"

"||v||=\sqrt{(\sqrt{2})^2+(\sqrt{2})^2}=\sqrt{4}=2"